multiple-choice test contains 10 questions. There are four possible answers for each question. a) In how many ways can a student answer the questions on the test if the student answers every question? b) In how many ways can a student answer the questions on the test if the student can leave answers blank?


1
Expert's answer
2021-12-24T10:08:04-0500

a) There are 4 possible answers for each question and we have 10 questions.

We can use the product rule: a student can answer the questions on the test in "4\times 4\times…\times 4=4^{10}" ways.

b) In this case there are 5 possible answers for each question and we have 10 questions.

In this case we can also use the product rule: a student can answer the questions on the test in "5\times 5\times…\times 5=5^{10}" ways.


Answers: a) "4^{10}"; b) "5^{10}" .